OpenStudy (anonymous):

I have a question how would you go about integrating S (x^2-1)/(x^(3/2))dx thank you

OpenStudy (anonymous):

As there is no product rule or quotient rule perform the division of each term and then integrate each one separately for example S(x^2/x^1.5) and S(-1//x^1.5)

OpenStudy (anonymous):

Simplify the exponents in the first term and then integrate like\[\int\limits x^n dx = \frac{x^{n+1}}{n+1}\]